Would someone be able to solve this question using the engineering economics equations? Mr.John purchases a car for 25,000 and finances his purchase by borrowing the money at 8% per year, compounded monthly. He pays off the loan with equal monthly payments for 5 years. What will the size of his monthly loan payment?
Answer #1

Initial Cost=Loan amount=L=$25000

Monthly rate of interest=i=8%/12=0.006667

Number of payments=n=5*12=60 (monthly)

Size of monthly payment=L*(A/P,i,n)=25000*(A/P,0.006667,60)

We need to calculate (A/P,0.006667,60)

Monthly payment=25000*0.020276=$506.90